OpenStudy (anonymous):

Please help!!! 3x^5 sqrt x + 7/x^2 sqrt x

OpenStudy (phi):

Can you add some parentheses. Not sure what you are saying. I assume you want to simplify this expression?

OpenStudy (anonymous):

its derivatives.

OpenStudy (anonymous):

(3x^2)(sqrt x) + 7/(x^2)(sqrt x)

OpenStudy (phi):

in the second term, what is in the denominator? \[x^{2} \sqrt{x}+\frac{7}{x^2 \sqrt{x}}\] ?

OpenStudy (anonymous):

7 divided by x sqaured times the square root of x

OpenStudy (phi):

One way to do this, is change sqrt(x) to x^(1/2) and add exponents for the second term, make it a negative exponent, and bring it up top.

OpenStudy (phi):

make sense?

OpenStudy (anonymous):

not really. ill be in tutoring today. I just had to figure this problem out. all my answers keep coming up wrong.

OpenStudy (phi):

You know \( x^{\frac{1}{2}} = \sqrt{x} \) and \( 3x^{5} \cdot x^{\frac{1}{2}} = x^{\frac{11}{2}} \) and for the second term you get \( 7x^{-\frac{5}{2}} \) assuming I can add fractions!

OpenStudy (phi):

The use the power rule \[\frac{d}{dx} x^{n}= n x^{n-1}\]

OpenStudy (phi):

* fixing typo: 3 x^{11/2} +7x^{-5/2}